OpenStudy (anonymous):

Which of the following expressions is the inverse of the function y = x - 2 / 3? y = 3x + 2 y = 2x + 3 y = x - 3 / by 2 y = x + 2 / by 3

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Is the original \[\Large y = x - \frac{2}{3}\] or is the original \[\Large y = \frac{x-2}{3}\]

OpenStudy (anonymous):

The second one

OpenStudy (usukidoll):

lack of sleep .. T_T

jimthompson5910 (jim_thompson5910):

\[\Large y = \frac{x-2}{3}\] swap x and y \[\Large x = \frac{y-2}{3}\] now solve for y

OpenStudy (usukidoll):

hmmm I would get rid of the denominator by multiplying 3 all the way first and then swap x and y.

jimthompson5910 (jim_thompson5910):

either way works

OpenStudy (anonymous):

I'm not sure what to start with when solving for y.

OpenStudy (usukidoll):

\[\Large 3y = (3)\frac{x-2}{3} \rightarrow 3y=x-2\]

OpenStudy (usukidoll):

then swap x and y . In other words switch places ...

OpenStudy (usukidoll):

\[\large 3x=y-2\]

OpenStudy (usukidoll):

and then solve for y...this would be much easier because I got rid of the fraction first by multiplying the entire equation by 3

OpenStudy (usukidoll):

so what do we need to do to have y by itself

OpenStudy (anonymous):

3x=y-2 +2 +2 5x=y

jimthompson5910 (jim_thompson5910):

3x and 2 are NOT like terms

OpenStudy (usukidoll):

ummmm 3x+2 is not combinable

jimthompson5910 (jim_thompson5910):

you cannot say 3x+2 turns into 5x or 5

OpenStudy (anonymous):

Oh. Sorry.

OpenStudy (usukidoll):

because 3x has a number and variable 2 is just a number... it can't be combined

OpenStudy (anonymous):

Ok... how do I isolate y.

OpenStudy (usukidoll):

y has to be by itself... \[\large 3x=y-2 \] so what do we need to do to have y by itself... y is destined to be alone xD

jimthompson5910 (jim_thompson5910):

unheard, when you added 2 to both sides, you were on the right path

jimthompson5910 (jim_thompson5910):

3x = y - 2 +2 +2 3x + 2 = y y = 3x + 2

OpenStudy (usukidoll):

:( ! why jim why

jimthompson5910 (jim_thompson5910):

you just leave 3x+2 as it is since you cannot combine those terms

OpenStudy (anonymous):

Oh, thank you Jim.

jimthompson5910 (jim_thompson5910):

she already had it usukidoll

OpenStudy (anonymous):

Thank you both.

OpenStudy (usukidoll):

jim's version would have required fractions to be split up \[\Large x = \frac{y-2}{3} \rightarrow x = \frac{y}{3}-\frac{2}{3}\] \[\Large x + \frac{2}{3} = \frac{y}{3}\] and then multiply 3 throughout the equation... too long...but another way to check \[ \Large 3x+2=y\]